OpenStudy (anonymous):

Solve the equation, using the domain of x = {-1, 0, 2}. 2x + 3y = 7

Parth (parthkohli):

Domain \(\Longrightarrow\) inputs. So we just have to solve for y. \( \color{Black}{\Rightarrow 2(-1) + 3y = 7}\) \( \color{Black}{\Rightarrow 2(0) + 3y = 7}\) \( \color{Black}{\Rightarrow 2(2) + 3y = 7 }\) Solve for y in all the three.